20190508-Analyzer-Mtg

From PREX Wiki
Jump to: navigation, search

Back to Main Page >> Analyzer_Meeting

previous meeting << >> following meeting

Logistic information

Meeting room at JLab: CC L210A (entrance is in the lobby across from the 2nd floor auditorium entrance)

 BlueJeans calling instructions:
 Toll-Free Number (U.S.&  Canada):  888-240-2560
 International toll number:         408-740-7256
 Bluejeans CODE:                  419 045 306
 Bluejeans link: https://bluejeans.com/419045306

Agenda

Analyzer code updates

  1. Recent branch activity & pull requests
    • PK: Tao's pull request for a different saturation level was merged. WD also had a pull to allow the data directory to be specified at the command line.
    • PK: two outstanding pull requests: a) Cameron's ellipticity calculation (almost ready). b) change to ET buffer. This needs to be tested before wrt performance.
  2. regression analysis
    • RR: japan seemed to be using the DeviceErrorFlag only (by comparing output fits to the regression output). This is being changed and at the same time we'll continue to compare a different run.
  3. feedback
    • CG: Amali just needs to implement the IA and helMag feedback.
    • PK: there is some strange lag at startup. Suspect we still have some hooks for old variables; will investigate. CG: will talk to Amali about it
  4. Beam modulation
    • Updates to the bmwClient
      • PK: will follow up later
    • Decoding and BMOD tree extraction
      • PK: will follow up later
  5. Others:
    • PK: we should clean up the rest of the branches.
    • PK: will put together a list of questionable issues.

Hardware/DAQ tasklist

  1. List of spares:
    • CC: 5 VQWKs (+2 in test stands); 2 FLEXIOs (we already have one in each crate -- tested). 5 HAPPEX timing boards (+5 more installed + 2 more in test crates). 4 spare VXworks ROCs.
    • CG: spare 7200? CC: found and tested (also in CH cabinite). PK: belongs to hall C and is generally used for Moller in Hall C.
  2. EPICS run-start and fast logger (and EPICS decoder to slow tree)
    • SP: we added the script to the msql DB and it seems to work now. There were some problems initially but not clear what solved them.
    • SP: the end of Run it runs a "end clean" and it kills all of these processes. The fastLogger gets cleaned properly. This should be included in the kcoda. PK: at least the fastLogger is included there.
    • CG: will compile list
  3. Timing: mostly understood, still small mysteries (LNE)
    • PK: to do tune beam scan we need to change how the VQWKs work (to make them work with the timing boards)
    • CC: installed but needs to be checked out. PK: until then we won't touch the gating. We'll work on it early next week.
  4. Synchronization-check system. w/help from Ed Jastrzembski
    • CC: haven't talked to him recently. Fiber cables will come in soon. Will check with him today.
    • PK: how about the old style? CC: plugged into injector but running into firmware with the scaler. PK: will look at it soon.
  5. Clean up cabling, labels, fasteners - First need to examine all signals
    • CC: in progress. Labeling will be done once timing system is fixed.
  6. Clean up CRL -- eliminate so-called "dirty trick"
    • CC: that is Bob’s court. PK: we should look at the CH and injector CRLs once we finish up with the timing and beamMod client.
  7. Pedestal studies, aka cross-talk
    • CC: KK talked to a group of us about pickup and we made a plan on how to intentionally putting in pickup noise and measure it and removing it to ppb levels.
    • CC: will write it down in a log entries.
  8. Green Monster rework: Adam.
    • AZ: made some progress. Will try to put it online.
  9. automatic production of run-dependent database
    • CG: still waiting on the computing center
  10. alarms like end-of-run voice
    • later
  11. e-logbook entries
    • later
  12. Hall C BCM signal
    • CC: plugged into scalar and checked that the cable is ok. Using BCM1 and 2.
  13. Document everything. Circuit diagrams, cable map, "how to".
    • CC: will make circuit diagrams.
  14. Talk to Brad about the system monitoring (HDD, CPU usage)
    • CG: will do that.
  15. Remote power supply cycling capability:
    • CC: will follow up with Alex C and Bob M.

ReSpin of APEX runs

runlist

    • What output do we have?
      • CC: 30 uA runs. HAPLOG entry 3755. CC: looked at the data and the 30 uA show similar issues with the 40/50 uA runs.
      • TY: will add respin scripts to the rootScripts folder.
      • SR: will add aggregator script the rootScripts folder.
      • CC: might be interesting to do a 2D residual plot from a Normal distribution.
      • CC: might be useful to have fixed scales for the 2D correlation plots.
    • standard plots
      • TY: working on the new plots.
  1. "Operations" branch
    • being updated.

Present

Sanghwa P, Cameron C, Tao Y, Paul K, Ciprian G, Devi A, Sakib R, Robert R, Wouter D

Excused

Kent P,